OpenStudy (anonymous):

A laser beam is incident on two slits with a separation of 0.225 mm, and a screen is placed 4.80 m from the slits. An interference pattern appears on the screen. If the angle from the center fringe to the first bright fringe to the side is 0.176°, what is the wavelength of the laser light?

OpenStudy (anonymous):

x=distance between center and bright spot=tan(.176)*4.8=.0147m d=Slit width=.225mm=0.000225m n=1because its the first bright spot l=distance between slits and screen=4.8m wavelength=xd/nl=6.89*10^-7m